OpenStudy (anonymous):

i raised to an even power can not be simplified to be -1, 1, i?

OpenStudy (anonymous):

i

OpenStudy (anonymous):

Thank you(:

OpenStudy (anonymous):

i^3 = i*i^2 i^3 = -i